Please answer this ASAP: Consider the energy market where supply is described by P = 25+5Q and demand is described by P=250-10Q. If the market is perfectly competitive and in equilibrium:

P* = $300 per unit and Q* = 55 units.

P* = $100 per unit and Q* = 15 units.

P* = $250 per unit and Q* = 45 units.

P*= $116.67 per unit and Q* = 18.33.

None of the above.
Answer #1

At equilibrium, demand is equal to supply.

From the supply equation,

Qs = (P - 25 )/5

Qd = (250 - P)/ 10

Qs = Qd

(P - 25)/5 = (250 -P)/10

P - 25 = (250 - P)/2

2P - 50 = 250 - P

3P = 300

P = 100

Qs = Qd = 15

P* = 100, Q* = 15